[figure: 9.7 cm]

MAGNETIC INVESTIGATIONS OF HUMAN LIVER TISSUES

E. Mosiniewicz-Szablewska^{a}, A. Ślawska-Waniewska^{a}, K. Światek^{a}, J. Galazka-Friedman^{b}, A. Friedman^{c}

aInstitute of Physics, Polish Academy of Sciences, Al. Lotnikow 32, 02-668 Warsaw,

Poland

bInstitute of Physics, Warsaw University of Technology, Koszykowa 75, Warsaw, Poland cDepartment of Neuorology, Warsaw Medical Academy, Banacha 1a, Warsaw, Poland

The lyophilized samples of the human liver have been investigated by means of electron paramagnetic resonance (EPR) spectroscopy and conventional magnetic methods over the temperature range 4 - 300 K. The EPR spectrum recorded at room temperature consists of a broad signal originating from the Fe^{3+} ions of ferritin grains. The behavior of this signal upon changes in temperature is characteristic of a system of particles that undergoes transition into a superparamagnetic state. By lowering the temperature the intensity of this signal decreases and below 8 K this line disappears, suggesting the formation of an antiferromagnetically ordered phase. Additionally, below 100 K the signal originating from Fe^{2+} ions of hemoglobin can be observed which shows a ferromagnetic-like behavior upon decrease in temperature. The temperature decay of d.c. susceptibility as well as ZFC-FC magnetization dependencies exhibit, besides of the diamagnetic contribution of the normal human tissue, two main magnetic contributions. One of them can be attributed to Fe^{2+} ions of hemoglobin and the second one, originating from the ferritin grains, which shows a transition into the superparamagnetic state at T>20 K.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : Dr Ewa Mosiniewicz-Szablewska

Address for correspondence : Institute of Physics, Polish Academy of Sciences, Al. Lotników 32/46 02-668 Warsaw, Poland

Email address : mosin@ifpan.edu.pl